OpenStudy (erniellyd28):

a moon is about 120,000 miles from the planet it orbits. a) write this number in scientific notation ( which i already did :-) ) b) if a rocket traveled 3*10 to the fourth power miles per hour, how long would it take for it to reach the moon from the planet?

OpenStudy (welshfella):

1.2 * 10^5 / 3 * 10^4

OpenStudy (welshfella):

you might find it easier to convert to standard numbers 120,000 and standard form of 3 * 10^4

OpenStudy (erniellyd28):

do i multiply the 2 answers to get my final answer?

OpenStudy (welshfella):

No you divide v = d / t so t = d / v

OpenStudy (erniellyd28):

but dont i suppose to multiply 3 times 10 to the fourth power?

OpenStudy (welshfella):

oh yes - you mean converting to standard form 3 * 10^4 = what in standard form?

OpenStudy (erniellyd28):

30,000

OpenStudy (erniellyd28):

am i correct?

OpenStudy (erniellyd28):

is the final answer 4 hours :-) ?

OpenStudy (welshfella):

yes 120,000 / 30,000 = 4 hours

OpenStudy (erniellyd28):

thank you very much :-)

OpenStudy (welshfella):

yw